In order to purchase a truck Shandex Transport (ST) obtained a $50 000 loan for 5 years at 7.8% compounded semi-annually. The loan is structured to reduce the balance owing to $10 000 at the end of the 5 year period. How much are ST's payments at the end of each month?
Answer #1

Using financial calculator
PV=-50000
FV=10000
N=12*5
I/Y=(1+7.8%/2)^(2/12)-1
CPT PMT=868.84

Payments at the end of each month=868.84